# Hierarchical Mamba Meets Hyperbolic Geometry: A New Paradigm for Structured Language Embeddings

### v2

@@ Abstract
Selective state-space models have achieved great success in long-sequence modeling. However, their capacity for language representation – especially in complex hierarchical reasoning tasks – remains underexplored. Most large language models rely on flat Euclidean embeddings, limiting their ability to capture latent hierarchies. To address this limitation, we propose Hierarchical Mamba (HiM), integrating efficient Mamba2 with exponential growth and curved nature of hyperbolic geometry to learn hierarchy-aware language embeddings for deeper linguistic understanding. Mamba2-processed sequences are projected to the Poincaré ball (via tangent-based mapping) or Lorentzian manifold (via cosine/sine-based mapping) with “learnable” curvature, optimized with a combined hyperbolic loss. Our HiM model facilitates the capture of relational distances across varying hierarchical levels, enabling effective long-range reasoning. This makes it well-suited for tasks like mixed-hop prediction and multi-hop inference in hierarchical classification. We evaluated our HiM with four linguistic and medical datasets for mixed-hop prediction and multi-hop inference tasks. Experimental results demonstrated that: 1) Both HiM models effectively capture hierarchical relationships for four ontological datasets, surpassing Euclidean baselines. 2) HiM-Poincaré captures fine-grained semantic distinctions with higher h-norms, while HiM-Lorentz provides more stable, compact, and hierarchy-preserving embeddings—favoring robustness over detail [^1] The source code is publicly available at https://github.com/BerryByte/HiM..
@@ 1 Introduction
Large language models (LLMs), such as Transformers [1] and BERT [2], typically encode input sequences into a flat Euclidean space. However, they struggle to capture the hierarchical and tree-like structures inherent in natural language [3], often leading to distortions at different levels of abstraction and specificity [4, 5]. Moreover, transformer-based encoders face significant computational overhead due to the quadratic complexity of the attention mechanism [1]. This limitation becomes particularly evident when dealing with hierarchical data (e.g., text ontologies, brain connectome [6, 7]) with exponentially expanding structure. State-space models, starting with the Structured State Space (S4) model [8], have shown exceptional scalability for long-sequence modeling. Mamba’s selective mechanism [9] dynamically prioritizes relevant information, achieving state-of-the-art performance in tasks with long-range dependencies. Mamba2 refines the original Mamba model for long-range sequence tasks by introducing a duality between state-space computations and attention-like operations, enabling the model to function as either an SSM or a structured, “mask-free” form of attention [10].
@@
Recently, leveraging hyperbolic geometry as the latent representation space in machine learning models has shown great promise for learning meaningful hierarchical structures [11, 12, 13]. The Poincaré disk and Lorentz model are two prevalent representations of hyperbolic space. The Poincaré disk model is often favored for its conceptual simplicity (bounded in a unit ball). However, the Lorentz model (with unbounded infinite space) offers a closed-form distance function, but requires careful handling of numerical functions dealing with space-like dimensions and a time-like dimension using exponential mapping and logarithmic mapping [12]. These numerical considerations are critical because the improper handling of the time-like coordinate in Lorentz models can lead to manifold violations, requiring specialized projection techniques [14, 15]. Existing hyperbolic LLM architectures [16, 12] often rely on Transformer blocks and apply a dimensionssimple Poincaré disk model, leading to $O(L^{2})$ complexity that becomes prohibitive for long sequences typical in deep hierarchies. A key challenge in implementing hyperbolic models is tuning the curvature to maintain numerical stability, particularly in Lorentz parameterization.
@@
In this paper, we introduce hyperbolic mamba with the Lorentz model, and compare it with its counterparts – Poincaré model and Euclidean model. To address the potential numerical instability in the Lorentz model [17], we explicitly bound the embedding norms and employ curvature-constrained Maclaurin approximations for hyperbolic operations. HiM aims to achieve high-performance hierarchical classification by preserving relational hierarchies. It demonstrates scalability for processing long sequences without compromising on accuracy or computational efficiency. HiM’s novelty lies in integrating a state-space model (Mamba2) with hyperbolic geometry, leveraging Mamba2’s $O(L)$ complexity for efficient sequence modeling while preserving hyperbolic properties for hierarchical representation. Additionally, our HiM incorporates novel task-specific hyperbolic losses that explicitly enforce parent-child distance constraints in hyperbolic space, enabling end-to-end hierarchy learning without Euclidean biases and achieving significant F1 gains on multi-hop inference tasks. To support HiM’s framework, we introduce SentenceMamba-16M, a compact, Mamba2-based large language model with 16 million parameters designed to generate high-quality sentence embeddings.
@@ 2 Related Works
Hyperbolic geometry has demonstrated strong potential in modeling hierarchical structures in both shallow and deep neural networks. Foundational works, such as Poincaré embeddings [4] and hyperbolic entailment cones [5], showed their effectiveness in capturing hierarchical relationships in taxonomies with shallow neural networks. Moreover, hyperbolic manifolds have also been applied to encode hierarchies in graph-structured data [18, 19]. More recent efforts have extended hyperbolic representations to multimodal computer vision tasks, including visual and audio modalities [20, 21], further demonstrating their strength in capturing both hierarchical structure and uncertainty.
@@
However, hyperbolic approaches in language modeling remain limited, although recent work has begun to extend them to transformers and their variants [16, 22, 23]. These approaches enable effective prediction of subsumption relations and transitive inferences across hierarchy levels using hyperbolic embeddings, providing a principled framework for encoding syntactic dependencies through geodesic distances. However, Hyperbolic BERT runs approximately 1.3$\times$ slower than standard BERT due to the complexity of hyperbolic operations [23]. To improve efficiency, recent works have explored fine-tuning LLMs directly in hyperbolic space with the Low-Rank Adaptation (LoRA) technique [24]. For example, HoRA [25] and HypLoRA [26] apply LoRA to the hyperbolic manifold, allowing parameter-efficient fine-tuning while capturing complex hierarchies. These methods show strong gains—up to 17.3% over Euclidean LoRA, but they mainly adopt a constant curvature, and may suffer from instability due to exponential/logarithmic mappings between Euclidean and hyperbolic spaces [27].
@@
Limitations in Current Approaches and Our Contribution: Despite significant progress, most existing methods either exploit only partial hyperbolic representations (e.g., using adapters or static embeddings) or rely heavily on attention-based architectures that scale poorly with long sequences and deep hierarchies. For instance, Poincaré GloVe [28] is limited to word embeddings, failing to capture dynamic, context-dependent relationships, while Hyperbolic BERT [23] and HiT [16] introduce significant computational overhead, especially for long sequences. Similarly, probing BERT’s embeddings in a Poincaré ball [22] to analyze hierarchical structures, but their diagnostic approach does not train a new model for hierarchical reasoning tasks like HiM. Methods, such as HoRA [25] and HypLoRA [26], only introduce hyperbolic geometry through adapter modules added post hoc to standard transformer backbones. These methods inherit the architectural inefficiencies of transformers but cannot fully encode hierarchy directly within the hyperbolic latent space.
Building on the strengths and limitations discussed above, we propose Hierarchical Mamba (HiM) as a novel framework for long-range hierarchical reasoning. Our contributions can be summarized as follows:
@@
  • SentenceMamba-16M: We introduce SentenceMamba-16M, a Mamba2-based LLM (16M parameters) trained at sentence-level embeddings on the SNLI dataset [29].
@@
  • Novel hyperbolic losses: HiM employs weighted clustering and centripetal losses to enforce parent-child separation and compact clustering of related entities with respect to origin, enhancing hierarchical structure preservation in hyperbolic space.
@@ 3 Methodology
3.1 Preliminaries
Hyperbolic geometry, characterized by negative curvature $\mathcal{K}=-1/c$, is well-suited for hierarchical data due to its exponential growth properties, modeled using the Poincaré ball or Lorentz model [4, 11]. Mamba2, a state-space model (SSM), offers efficient sequence modeling with linear complexity, using structured state-space duality to balance SSM and attention-like operations [10]. The detailed formulations and preliminaries of Mamba2 are provided in Appendix A.2.
@@ 3.2 Hyperbolic Mamba (HiM)
The overall framework of HiM, including the integration of Mamba2 blocks and hyperbolic projections, is shown in Fig. 1. Firstly, the raw text is tokenized into a sequence of tokens; these token IDs are mapped into embedded tokens resulting in token IDs of shape $[B,L,D,N]$, where $B$ is the batch size ($B=256$), $L$ is the sequence length ($L=128$), and $D$ is the embedding dimension ($D=384$), and N is the state dimension ($N=96$). These embeddings are passed through a sequence of 4 Mamba2 Blocks. Alanis-Lobato et al. [30] focus on efficient embedding of complex networks into hyperbolic space using the network Laplacian, achieving a computational complexity of $O(N^{2})$ and enabling the analysis of large networks in seconds. This highlights the importance of computational efficiency in scaling hyperbolic models, a principle that HiM extends by Mamba2 blocks (Equations 17 and 20) to achieve linear-time complexity $O(L)$, making it particularly suited for long sequences and deep hierarchical language structures. In the Mamba2 blocks, the inputs $B$ having 384 dimensions are projected into the intermediate state $I$ having 768 dimensions using a linear transformation ($D:384\to I:768$).
The $x_{t}^{\prime}$ component undergoes a convolution operation with a kernel size of 4. A SiLU activation function follows this operation for non-linearity. A detailed implementation of the Mamba2 block is discussed in Appendix A.2. In the output projection, the intermediate state dimensions are projected back to their original embedding dimension $I:768\to D:384$ for compatibility with downstream tasks. The SentenceMamba-16M model, central to HiM, is randomly initialized with Kaiming normal weights instead of pretrained weights, enabling it to learn hierarchical structures directly from training data in hyperbolic space without any biases. The SentenceMamba-16M model is trained using Triplet Contrastive Loss, which brings embeddings of positive pairs closer together while pushing apart embeddings of other sentences in the batch. This loss function has proven effective in prior works involving hierarchical embedding [31, 16]. After normalizing each embedding to unit length, we measure the pairwise cosine similarity as $\text{sim}(i,j)\;=\;\mathbf{e}_{i}\cdot\mathbf{e}_{j}$, where each embedding $\mathbf{e}_{i}$ and $\mathbf{e}_{j}$ belong to $\mathbf{e}\in\{\mathbf{e}_{1},\mathbf{e}_{2},\dots,\mathbf{e}_{n}\}$. We then calculate the contrastive loss for the batch by constructing a similarity matrix from the similarity scores across nodes. The sentence embeddings are constrained using hyperbolic tangent activation followed by L2 normalization to ensure numerical stability:
@@
This approach mitigates numerical overflow and enhances training stability. The interplay between the norm scaling parameter $\gamma$ and curvature $\mathcal{K}=-1/c$ is mathematically significant. When we scale embeddings by $\gamma$ before projection, it effectively modulates the curvature of the hyperbolic space. For the curvature parameter $c$, scaling the norm of $\mathbf{h}$ by $\gamma$ leads to an effective curvature of $\mathcal{K}_{\text{eff}}=-1/(c\gamma^{2})$. This can be seen from the Poincaré ball projection equation (3), where scaling $\|\mathbf{h}\|$ by $\gamma$ is equivalent to modifying $\sqrt{c}$ by a factor of $1/\gamma$. Thus, the model learns both a base curvature $c$ and a scaling factor $\gamma$ that together determine the optimal geometry for representing hierarchical relationships. This dual learning approach provides flexibility in adapting the hyperbolic space to the structural complexity of the data while maintaining numerical stability. Then the vector $h$ is mapped to a point $e$ in hyperbolic space. The general form for a Poincaré ball with radius $r=\sqrt{c}$ (curvature $\mathcal{K}=-1/c=-1/r^{2}$) is:
@@
where $\|h\|$ is the norm of $h$. This scaling ensures the vector lies within the unit ball. This yields the final sentence embedding $e$ with values constrained between $-1$ and $1$ (indicating a positive or negative parent), allowing us to project the embedding onto the Poincaré Ball manifold.
@@
Here, $h$ is the norm distance, $\sqrt{c}$ is the radius of the hyperbolic space; $\mathcal{K}<0$ ensures hyperbolic geometry. For the Lorentz mapping we let $z=\|\mathbf{h}\|/\sqrt{c}$ in Equation 4. The cosh, sinh functions are the Hyperbolic cosine and sine functions used to compute projections. The first term $\sqrt{c}\cdot\cosh\left(\frac{\|\mathbf{h}\|}{\sqrt{c}}\right)$ in the Lorentz projection is the time-like dimension. The remaining components $\sqrt{c}\cdot\sinh\left(\frac{\|\mathbf{h}\|}{\sqrt{c}}\right)\cdot\frac{\mathbf{h}}{\|\mathbf{h}\|}$ are the space-like dimension. This step is crucial for hyperbolic geometry as it ensures the embeddings are bounded, enabling seamless projection onto the Lorentzian manifold.
@@
[…]
@@
To fully exploit the hyperbolic structure of our model, we employ an advanced hyperbolic loss function for the HiM model optimization, which is a weighted combination of centripetal loss and clustering loss. These losses enhance the model’s ability to effectively learn hierarchical relationships by optimally positioning and grouping the embeddings in a strongly hierarchical structure within the Hyperbolic manifold. Detailed equations for our hyperbolic loss are presented below, with the full calculation process provided in Appendix B.
@@
where $w_{ce}$ and $w_{cl}$ are weights that control the contribution of each loss component. This loss ensures that the model maintains the hierarchical structure during training, with parent entities closer to the origin and related entities clustered together. The margins $\alpha$ and $\beta$ in the clustering and centripetal losses (Equations 7 and  8) are implemented as dynamic parameters optimized during training to adaptively enforce the hierarchical constraints. The margins for the clustering and centripetal losses are dependent on the radius ($r=\sqrt{c}$) that is dependent on the curvature $\mathcal{K}$, the margins are updated over the training steps in terms of the radius $r$ as: $\alpha=0.255\cdot r$; $\beta=0.0051\cdot r$. These scaling factors were determined through empirical validation to maintain consistent separation properties as the model adapts its curvature during training. The clustering margin is intentionally larger to enforce robust hierarchical separation between related and unrelated entities, while the centripetal margin is smaller to allow fine-grained positioning of parent nodes closer to the origin relative to their children, reflecting the natural expansion of hierarchies in hyperbolic space.
In all hierarchical classification tasks, hard negatives were chosen to sharpen the model’s discrimination [31]. Rather than randomly sampling unrelated nodes, we select negative examples that are semantically close to the anchor (or positive) in embedding space. This training strategy forces the model to learn more subtle hierarchical distinctions, which is crucial for tasks such as “multi-hop inference”. We observe that hard negatives lead to better generalization.
Following [19], we optimize the learnable curvature parameter using the AdamW optimizer. This is justified because the curvature parameter itself is a scalar Euclidean variable controlling the hyperbolic manifold geometry, making AdamW both theoretically valid and empirically stable. To ensure numerical stability during training in hyperbolic space as the curvature adapts, we implement a geometric stabilization technique that periodically projects the model parameters back onto the manifold. Specifically, every 100 optimization steps, this stabilization counteracts numerical drift that can occur during curvature optimization, preventing embeddings from violating the constraints of the hyperbolic geometry and ensuring all distance computations remain well-defined throughout training.

### v3

@@ Abstract
Selective state-space models excel at long-sequence modeling, but their capacity for language representation – in complex hierarchical reasoning – remains underexplored. Most large language models rely on flat Euclidean embeddings, limiting their ability to capture latent hierarchies. To address this, we propose Hierarchical Mamba (HiM), integrating efficient Mamba2 with hyperbolic geometry to learn hierarchy-aware language embeddings for deeper linguistic understanding. Mamba2-processed sequences are projected to the Poincaré ball or Lorentzian manifold with “learnable” curvature, optimized with a hyperbolic loss. Our HiM model facilitates the capture of relational distances across varying hierarchical levels, enabling effective long-range reasoning for tasks like mixed-hop prediction and multi-hop inference in hierarchical classification. Experimental results show both HiM variants effectively capture hierarchical relationships across four linguistic and medical datasets, surpassing Euclidean baselines, with HiM-Poincaré providing fine-grained distinctions with higher h-norms, while HiM-Lorentz offers more stable, compact, and hierarchy-preserving embeddings-favoring robustness [^1] The source code is publicly available at https://github.com/BerryByte/HiM.
@@ 1 Introduction
Large language models (LLMs), such as Transformers (Vaswani, 2017) and BERT (Devlin et al., 2019), typically encode input sequences into a flat Euclidean space. However, they struggle to capture the hierarchical and tree-like structures inherent in natural language (Chomsky, 2014), often leading to distortions at different levels of abstraction and specificity (Nickel & Kiela, 2017; Ganea et al., 2018). Moreover, transformer-based encoders face significant computational overhead due to the quadratic complexity of the attention mechanism (Vaswani, 2017). This limitation becomes particularly evident when dealing with hierarchical data (e.g., text ontologies, brain connectome (Ramirez et al., 2025; Baker et al., 2024)) with exponentially expanding structure. State-space models, starting with the Structured State Space (S4) model (Gu et al., 2021), have shown exceptional scalability for long-sequence modeling. Mamba’s selective mechanism (Gu & Dao, 2023) dynamically prioritizes relevant information, achieving state-of-the-art performance in tasks with long-range dependencies. Mamba2 refines the original Mamba model for long-range sequence tasks by introducing a duality between state-space computations and attention-like operations, enabling the model to function as either an SSM or a structured, “mask-free” form of attention (Dao & Gu, 2024).
@@
Recently, leveraging hyperbolic geometry as the latent representation space in machine learning models has shown great promise for learning meaningful hierarchical structures (Nickel & Kiela, 2018; Peng et al., 2021; Petrovski, 2024). The Poincaré disk and Lorentz model are two prevalent representations of hyperbolic space. The Poincaré disk model is often favored for its conceptual simplicity (bounded in a unit ball). However, the Lorentz model (with unbounded infinite space) offers a closed-form distance function, but requires careful handling of numerical functions dealing with space-like dimensions and a time-like dimension using exponential mapping and logarithmic mapping (Peng et al., 2021). These numerical considerations are critical because the improper handling of the time-like coordinate in Lorentz models can lead to manifold violations, requiring specialized projection techniques (Fan et al., 2024; Liang et al., 2024). Existing hyperbolic LLM architectures (He et al., 2024b; Peng et al., 2021) often rely on Transformer blocks and apply a simple Poincaré disk model, leading to $O(L^{2})$ complexity that becomes prohibitive for long sequences typical in deep hierarchies. A key challenge in implementing hyperbolic models is tuning the curvature to maintain numerical stability, particularly in Lorentz parameterization.
@@
In this paper, we introduce hyperbolic mamba with the Lorentz model, and compare it with its counterparts – Poincaré model and Euclidean model. To address the potential numerical instability in the Lorentz model (Mishne et al., 2023), we explicitly bound the embedding norms and employ curvature-constrained Maclaurin approximations for hyperbolic operations. HiM aims to achieve high-performance hierarchical classification by preserving relational hierarchies. It demonstrates scalability for processing long sequences without compromising on accuracy or computational efficiency. HiM’s novelty lies in integrating a state-space model (Mamba2) with hyperbolic geometry, leveraging Mamba2’s $O(L)$ complexity for efficient sequence modeling while preserving hyperbolic properties for hierarchical representation. Additionally, our HiM incorporates task-specific hyperbolic losses that explicitly enforce parent-child distance constraints in hyperbolic space, enabling end-to-end hierarchy learning without Euclidean biases and achieving significant F1 gains on multi-hop inference tasks. To support HiM’s framework, we introduce SentenceMamba-16M, a compact, Mamba2-based large language model with 16 million parameters designed to generate high-quality sentence embeddings.
@@ 2 Related Works
Hyperbolic geometry has demonstrated strong potential in modeling hierarchical structures in both shallow and deep neural networks. Foundational works, such as Poincaré embeddings (Nickel & Kiela, 2017) and hyperbolic entailment cones (Ganea et al., 2018), showed their effectiveness in capturing hierarchical relationships in taxonomies with shallow neural networks. Moreover, hyperbolic manifolds have also been applied to encode hierarchies in graph-structured data (Liu et al., 2019; Chami et al., 2019). More recent efforts have extended hyperbolic representations to multimodal computer vision tasks, including visual and audio modalities (Yang et al., 2024c; Mandica et al., 2024), further demonstrating their strength in capturing both hierarchical structure and uncertainty.
@@
However, hyperbolic approaches in language modeling remain limited. As an early approach to hyperbolic word embeddings, Dhingra et al. (2018) provided an important early step by reparameterizing Poincaré embeddings for GRU-based sequence modeling. This approach eliminated projection steps and supported both shallow and parametric encoders, but it was ultimately limited by its shallow representations, which restricted its expressive power and ability to capture long-range dependencies. More recent research has extended these concepts to transformers and their variants (He et al., 2024b; Chen et al., 2021; 2024). These approaches enable effective prediction of subsumption relations and transitive inferences across hierarchy levels using hyperbolic embeddings, providing a principled framework for encoding syntactic dependencies through geodesic distances. However, Hyperbolic BERT exhibits high computational cost than standard BERT due to the complexity of hyperbolic operations (Chen et al., 2024). To improve efficiency, recent works have explored fine-tuning LLMs directly in hyperbolic space with the Low-Rank Adaptation (LoRA) technique (Hu et al., 2022). For example, HoRA (Yang et al., 2024a) and HypLoRA (Yang et al., 2024b) apply LoRA to the hyperbolic manifold, allowing parameter-efficient fine-tuning while capturing complex hierarchies. These methods show strong gains—up to 17.3% over Euclidean LoRA. However, these models usually assume a constant curvature, which may not be optimal for all data, and can suffer from numerical instability due to the exponential and logarithmic mappings required to transition between Euclidean and hyperbolic spaces (López & Strube, 2020; Patil et al., 2025).
@@
Limitations in Current Approaches and Our Contribution: Despite significant progress, most existing methods either exploit only partial hyperbolic representations (e.g., using adapters or static embeddings) or rely heavily on attention-based architectures that scale poorly with long sequences and deep hierarchies. For instance, Poincaré GloVe (Tifrea et al., 2018) is limited to word embeddings, failing to capture dynamic, context-dependent relationships, while Hyperbolic BERT (Chen et al., 2024) and HiT (He et al., 2024b) introduce significant computational overhead, especially for long sequences. Similarly, probing BERT’s embeddings in a Poincaré ball (Chen et al., 2021) to analyze hierarchical structures, but their diagnostic approach does not train a new model for hierarchical reasoning tasks like HiM. Methods, such as HoRA (Yang et al., 2024a) and HypLoRA (Yang et al., 2024b), only introduce hyperbolic geometry through adapter modules added post hoc to standard transformer backbones. These methods inherit the architectural inefficiencies of transformers but cannot fully encode hierarchy directly within the hyperbolic latent space. Building on the strengths and limitations discussed above, we propose Hierarchical Mamba (HiM) as a novel framework for long-range hierarchical reasoning. Our contributions can be summarized as follows:
@@
  • SentenceMamba-16M: We introduce SentenceMamba-16M, a Mamba2-based LLM (16M parameters) trained at sentence-level embeddings on the SNLI dataset (Bowman et al., 2015).
@@
  • Hyperbolic losses: HiM employs weighted clustering and centripetal losses to enforce parent-child separation and compact clustering of related entities with respect to origin, enhancing hierarchical structure preservation in hyperbolic space.
@@ 3 Methodology
Hyperbolic geometry, characterized by negative curvature $\mathcal{K}=-1/c$, is well-suited for hierarchical data due to its exponential growth properties, modeled using the Poincaré ball or Lorentz model (Nickel & Kiela, 2017; 2018). Mamba2, a state-space model (SSM), offers efficient sequence modeling with linear complexity, using structured state-space duality to balance SSM and attention-like operations (Dao & Gu, 2024). The detailed formulations and preliminaries of Mamba2 are provided in Appendix A.1.
@@ 3.1 Hyperbolic Mamba (HiM)
The overall framework of HiM, including the integration of Mamba2 blocks and hyperbolic projections, is shown in Figure 1. Firstly, the raw text is tokenized into a sequence of tokens; these token IDs are mapped into embedded tokens resulting in token IDs of shape $[B,L,D,N]$, where $B$ is the batch size ($B=256$), $L$ is the sequence length ($L=128$), and $D$ is the embedding dimension ($D=384$), and N is the state dimension ($N=96$). These embeddings are passed through a sequence of 4 Mamba2 Blocks. Alanis-Lobato et al. (2016) focuses on efficient embedding of complex networks into hyperbolic space using the network Laplacian, achieving a computational complexity of $O(N^{2})$ and enabling the analysis of large networks in seconds. This highlights the importance of computational efficiency in scaling hyperbolic models, a principle that HiM extends by Mamba2 blocks (Equations 17 and 20) to achieve linear-time complexity $O(L)$, making it particularly suited for long sequences and deep hierarchical language structures. In the Mamba2 blocks, the inputs $B$ having 384 dimensions are projected into the intermediate state $I$ having 768 dimensions using a linear transformation ($D:384\to I:768$).
@@
The $x_{t}^{\prime}$ component undergoes a convolution operation with a kernel size of 4. A SiLU activation function follows this operation for non-linearity. A detailed implementation of the Mamba2 block is discussed in Appendix A.1. In the output projection, the intermediate state dimensions are projected back to their original embedding dimension $I:768\to D:384$ for compatibility with downstream tasks. The SentenceMamba-16M model, central to HiM, is randomly initialized with Kaiming normal weights instead of pretrained weights, enabling it to learn hierarchical structures directly from training data in hyperbolic space without any biases. The SentenceMamba-16M model is trained using Triplet Contrastive Loss, which brings embeddings of positive pairs closer together while pushing apart embeddings of other sentences in the batch. This loss function has proven effective in prior works involving hierarchical embedding (Schroff et al., 2015; He et al., 2024b). After normalizing each embedding to unit length, we measure the pairwise cosine similarity as $\text{sim}(i,j)\;=\;\mathbf{e}_{i}\cdot\mathbf{e}_{j}$, where each embedding $\mathbf{e}_{i}$ and $\mathbf{e}_{j}$ belong to $\mathbf{e}\in\{\mathbf{e}_{1},\mathbf{e}_{2},\dots,\mathbf{e}_{n}\}$. We then calculate the contrastive loss for the batch by constructing a similarity matrix from the similarity scores across nodes. The sentence embeddings are constrained using hyperbolic tangent activation followed by L2 normalization to ensure numerical stability:
@@
This approach mitigates numerical overflow and enhances training stability. The interplay between the norm scaling parameter $\gamma$ and curvature $\mathcal{K}=-1/c$ is mathematically significant. When we scale embeddings by $\gamma$ before projection, it effectively modulates the spread of embeddings in the hyperbolic space. Our model learns both the curvature parameter $c$ and a scaling factor $\gamma$ that together determine the optimal geometry for representing hierarchical relationships. This dual learning approach provides flexibility in adapting the hyperbolic space to the structural complexity of the data while maintaining numerical stability. Then the vector $h$ is mapped to a point $e$ in hyperbolic space. The general form for a Poincaré ball with radius $r=\sqrt{c}$ (curvature $\mathcal{K}=-1/c=-1/r^{2}$) is:
@@
where $\|h\|$ is the norm of embedding vector $h$. This scaling ensures the vector lies within the unit ball. This yields the final sentence embedding $e$ with values constrained between 1-1 and $1$ (indicating a positive or negative parent), allowing us to project the embedding onto the Poincaré Ball manifold.
@@
Here, $\|h\|$ is the norm distance of embedding vector $h$, $\sqrt{c}$ is the radius of the hyperbolic space; $\mathcal{K}<0$ ensures hyperbolic geometry. For the Lorentz mapping we let $z=\|\mathbf{h}\|/\sqrt{c}$ in Equation 4. The cosh, sinh functions are the Hyperbolic cosine and sine functions used to compute projections. The first term $\sqrt{c}\cdot\cosh\left(\frac{\|\mathbf{h}\|}{\sqrt{c}}\right)$ in the Lorentz projection is the time-like dimension. The remaining components $\sqrt{c}\cdot\sinh\left(\frac{\|\mathbf{h}\|}{\sqrt{c}}\right)\cdot\frac{\mathbf{h}}{\|\mathbf{h}\|}$ are the space-like dimension. This step is crucial for hyperbolic geometry as it ensures the embeddings are bounded, enabling seamless projection onto the Lorentzian manifold.
@@
$\cosh(z)\;=\;1\;+\;\frac{z^{2}}{2!}\;+\;\frac{z^{4}}{4!}\;+\;\cdots;\quad\quad\sinh(z)\;=\;z\;+\;\frac{z^{3}}{3!}\;+\;\frac{z^{5}}{5!}\;+\;\cdots.$ (5)
@@
To fully exploit the hyperbolic structure of our model, we employ an advanced hyperbolic loss function for the HiM model optimization, which is a weighted combination of centripetal loss and clustering loss. These losses enhance the model’s ability to effectively learn hierarchical relationships by optimally positioning and grouping the embeddings in a strongly hierarchical structure within the Hyperbolic manifold. Detailed equations for our hyperbolic loss are presented below, with the full calculation process provided in Appendix A.2.
@@
where $w_{ce}$ and $w_{cl}$ are weights that control the contribution of each loss component. This loss ensures that the model maintains the hierarchical structure during training, with parent entities closer to the origin and related entities clustered together. The margins $\alpha$ and $\beta$ in the clustering and centripetal losses (Equations 7 and  8) are implemented as dynamic parameters optimized during training to adaptively enforce the hierarchical constraints. The margins for the clustering and centripetal losses are adapted to the hyperbolic geometry by scaling proportionally with the radius $r=\sqrt{c}$, ensuring the loss functions remain geometrically consistent across different curvatures. These scaling factors were determined through empirical validation to maintain consistent separation properties as the model adapts its curvature during training. The clustering margin is intentionally larger to enforce robust hierarchical separation between related and unrelated entities, while the centripetal margin is smaller to allow fine-grained positioning of parent nodes closer to the origin relative to their children, reflecting the natural expansion of hierarchies in hyperbolic space. In all hierarchical classification tasks, hard negatives were chosen to sharpen the model’s discrimination (Schroff et al., 2015). Rather than randomly sampling unrelated nodes, we select negative examples that are semantically close to the anchor (or positive) in embedding space. This training strategy forces the model to learn more subtle hierarchical distinctions, which is crucial for tasks such as “multi-hop inference”. We observe that hard negatives lead to better generalization.
Following Chami et al. (2019), we optimize the learnable curvature parameter using the AdamW optimizer. This is justified because the curvature parameter itself is a scalar Euclidean variable controlling the hyperbolic manifold geometry, making AdamW both theoretically valid and empirically stable. To ensure numerical stability during training in hyperbolic space as the curvature adapts, we implement a geometric stabilization technique that periodically projects the model parameters back onto the manifold. Specifically, every 100 optimization steps, this stabilization counteracts numerical drift that can occur during curvature optimization, preventing embeddings from violating the constraints of the hyperbolic geometry and ensuring all distance computations remain well-defined throughout training.